navy seal carries four types of weapons to carry a secret mission alone. Assume that the probability for cach weapon to fail to operate is 1%. Calculate the probability belovw 0 points] A (a) All four weapons cannot operate properly. (b) At least one weapon works.
Answer #1

Here we have; P o.DI q = I-P= 1-o. 이 0.99 ㄐㄧㄚ O Phey): чсч ь.oiro.ggy-y At least one weapon workr 0 10.9606